OpenStudy (anonymous):

if 2t=e^x-1 then x=?

OpenStudy (psymon):

Well, isolate the e^x expression first. 2t+1 = e^x. Now normally when you're solving for x you do opposites, right? If you have a + 5 then you subtract 5 from both sides. If you have 2x, you divide both sides by 2. So that being said, what would be the opposite of e^x?

OpenStudy (anonymous):

2t=e^(x-1)

OpenStudy (psymon):

Ah, I apologize. Well then your e^(x-1) expression is already isolated. Now you would just do the inverse of e^x to both sides of the equation.

OpenStudy (anonymous):

ln(2t)+1=x ?

OpenStudy (psymon):

Yep :3

OpenStudy (anonymous):

Thank you!!!

OpenStudy (psymon):

Of course : )